# LAG: LLM agents for Leaderboard Auto Generation on Demanding

### v1

@@ 4.2.3 Meta Evaluation (RQ-3)
Figure 4: Pearson Correlation Coefficient values given by four LLMs and human experts. Note that the Pearson Correlation Coefficient is between -1 and 1, the larger value indicates more positive correlations.
To verify the consistency between our proposed LLM evaluation strategy and human evaluation, we conduct a correlation evaluation involving human experts and our automated evaluation method. Human experts judge pairs of generated leaderboards to determine which one is superior. We compare the judgments made by our method against those made by human experts. Specifically, we provide experts with the same scoring criteria used in our evaluation for reference. The experts rank the 20 LAG-generated leaderboards, and we compare these rankings with those generated by the LLM using the Pearson Correlation Coefficient to measure consistency between human and LLM evaluations.
The results of this meta-evaluation are presented in Figure 4. The table shows the Pearson Correlation Coefficient values, indicating the degree of correlation between the rankings given by each LLM and the human experts. The Pearson correlation coefficient values indicate a strong positive correlation between the quality scores provided by the LLM and those given by human experts, with the O1-preview achieving the highest correlation at 0.76. These results suggest that our evaluation method aligns well with human preferences, providing a reliable proxy for human evaluation.
4.2.4 Ablation Study (RQ-4)
To understand the contribution of each component in LAG, we conducted an ablation study by removing key components of LAG as follows: (1) LAG w/o Table Classification: We removed the table classification step, which led to a slight decrease in Structure and Multiaspect scores, indicating that classifying tables is essential for maintaining a logical and well-organized leaderboard. (2) LAG w/o Refinement: We disabled the Refinementing step, which resulted in a minor drop in Coverage and Latest scores, suggesting that Refinementing helps refine the leaderboard by ensuring that only the most relevant and recent papers are included. As shown in Table 3, the results of the ablation study confirm that each component of the LAG plays a crucial role in achieving the generation of the high-quality leaderboard.
5 Conclusion
We introduce LAG, a novel agent framework leveraging large language models to automatically generate the latest, and high-quality leaderboards based on given research topics. LAG addresses key challenges including paper coverage, fair comparison, and timeliness through a systematic approach involving paper collection and split, table extraction and classification, table unpacking and integration, and leaderboard generation and evaluation. Experiments showed that LAG can automatically generate new, high-quality leaderboards in a relatively short time and match human performance topic-related quality and content qulaity. This advancement offers a scalable and effective solution for synthesizing the latest leaderboards, providing a valuable tool for researchers in rapidly evolving fields like artificial intelligence.
Limitations
One limitation of LAG is its reliance on the quality of the retrieved papers. While our topic-related quality metrics are strong, there is still room for improvement in ensuring that all relevant papers are included. Future work could explore more sophisticated retrieval models to further enhance the coverage of the generated leaderboards. Another limitation is, a specific dataset may contain several evaluation metrics, and different papers may use different metrics to evaluate proposed models’ performance, bringing challenges for leaderboard generation and baseline comparison.
Appendix A Example Prompts
The prompts of instructing LLMs in different stages of LAG are illustrated in Prompts 1, 2, and 3.
<instruction> You are an expert in summarizing and extracting key content from LaTeX-formatted academic papers on computers and artificial intelligence. Please output your reply in the following JSON format: <format>[EXAMPLE JSON]</format> ======================================================================== There are some key points to note: In the "selected table’s core results", other models’ results are of no concern and should be omitted. The table’s header metrics should be the same as the evaluation metrics chosen. The number of items in the "classification of tables" dict should be equal to the "number of tables" int value. These two items help you to identify the main result tables better. All three items about the settings in the JSON output should be corresponding to the proposed method’s best performance in the selected table. Sometimes in the selected table, the proposed method’s performance may not be unique (e.g., different hyperparameters or training strategies), you need to choose the best one and it usually appears in the last row of the table. If there are multiple tables that meet the requirements (both being the main result table and based on the specified dataset), choose the one with richer information. ======================================================================== Here, I provide you with an example of the complete process to help you understand your task. First, I provide you an article: <article>[EXAMPLE ARTICLE]</article> Afterwards, I specify the dataset as [EXAMPLE DATASET], you should output: <format>[EXAMPLE RESPONSE]</format> </instruction>
Box 1 The prompt of the table extraction agent, with the table classification COT procedure.
<instruction> You are an expert in summarizing and extracting key content from LaTeX-formatted academic papers on computers and artificial intelligence. Please output your reply in the following JSON format: <format>[EXAMPLE JSON]</format> ======================================================================== There are some key points to note: In the "main result table’s core results", other models’ results are of no concern and should be omitted. The main result table’s header metrics should be the same as the evaluation metrics chosen. All three items about the settings in the JSON output should correspond to the proposed method’s best performance in the selected table. Sometimes in the main result table, the proposed method’s performance may not be unique (e.g., different hyperparameters or training strategies), you need to choose the best one and it usually appears in the last row of the table. If there are multiple tables that meet the requirements (both being the main result table and based on the specified dataset), choose the one with richer information. ======================================================================== Here, I provide you with an example of the complete process to help you understand your task. First, I provide you an article: <article>[EXAMPLE ARTICLE]</article> Afterwards, I specify the dataset as [EXAMPLE DATASET], you should output: <format>[EXAMPLE RESPONSE]</format> </instruction>
Box 2 The prompt of the table extraction agent, w/o the table classification COT procedure.
<instruction> You are an expert in constructing the Artificial Intelligence leaderboard. Please refer to the content I provide you to answer the user’s questions. The contents I provide you are a number of structured summaries extracted from computer/artificial intelligence papers. You need to build a markdown format leaderboard (showcase the performance of the models on the same dataset, each line representing a specific model) based on the titles, experimental settings, and evaluation metrics of these articles. Please output your reply in the Markdown format. Here, I list a complete example of the question and the answer to help you understand your task. For example, I provide you a list of JSON files containing the extracted content of the articles: [JSON LIST] The expected leaderboard that you generate should be: [EXAMPLE LEADERBOARD] ======================================================================== Pay attention: The leaderboard should be in the Markdown format and reflect all the articles provided! The leaderboard in the dictionary format is forbidden! In the above case, selecting Pre and Rec as the metrics in the final leaderboard is not appropriate because in most articles the corresponding performance values are absent. ======================================================================== Here, the target list of extracted content of the articles is as follows: [TARGET JSON LIST] Please give me the well-organized leaderboard of the provided articles. The leaderboard should have consistent performance metrics, github links, settings and the title of the article, etc. The leaderboard should be in the Markdown format and reflect all the articles provided. Warning: I need a well-organized markdown-format leaderboard containing all the articles’ information. The leaderboard’s max serial number in the "No." column should equal to the number of articles provided. When selecting metrics, you need to consider their text descriptions. The same metric may have multiple different abbreviations. In the final table, there must not be any duplicate metrics (it is unacceptable to have duplicates where different abbreviations represent the same meaning). Large-scale omissions are not allowed! For each model, only a small portion of the results are missing under the selected metrics. The vast majority of the metrics have corresponding values. The abbreviations for the same metric may be different, but you need to avoid being misled by the abbreviations. Use approximate intersections to select metrics from the given articles, while avoiding a large amount of data waste. Allow some models to have a certain degree of data missing under the selected metrics. The content in the "Experimental Setting" column should be concise and non-descriptive, just a few words. When different articles use different units for the same metric, please note that you need to convert them when integrating them so that the units in the final leaderboard are consistent. For example, 50% is equal to 0.5. "50" and "0.5" should not be presented in the same column of a leaderboard. Check each column corresponding to the selected metrics in the final leaderboard. If more than 60% of the values in that column are missing or represented by placeholders, the metric should be discarded. </instruction>
Box 3 The prompt of the leaderboard construction agent.
1{
2 "title": "The title of the paper (String)",
3 "number of tables": "The number of tables in the paper, denoted as n (Int)",
4 "classification of tables": "The classification of tables in the paper, including 4 types: main result/comparison tables(0), ablation tables(1), hyperparameter tables(2), and others(3). You should output a dictionary with the number of tables and their corresponding types, formed as {"0":0, "1":2, ..., "n-1":3} (Dict)",
5 "selected table’s index": "The index of the main result table focused on the specified dataset [SPECIFIED DATASET], denoted as i (Int)",
6 "metrics": "The evaluation metrics chosen to assess performance of the method proposed in this paper. This information is extracted from the textual portion of the ’Experimental’ related section (String)",
7 "selected table’s metrics": "Metrics used in the selected main result table, it should be almost the same as the metrics extracted from the textual. Remove the latex format syntax (String)",
8 "selected table’s core results": "A dictionary only containing this paper’s model best performance on the selected dataset, with the metrics as keys and the corresponding values (Dict)",
9 "selected table’s settings (model & size)": "In computer vision, the model usually means the backbone architecture of the network, such as ResNet, ViT, and so on. The size can be omitted if not specified. In NLP, the model and size are usually organized as a string, such as ’LLAMA-7B’, ’GPT-3’, and so on (String)",
10 "selected table’s settings (training strategy)": "Training strategy usually refers to the concepts like: fine-tuning, transfer learning, linear-probing, reinforce learning, one-shot, few-shot, prompt-learning, semi/self supervised and so on (String)",
11 "selected table’s settings (hyperparameter selection)": "The hyperparameters used in the model, such as learning rate, batch size, and so on. You should output a dictionary with the hyperparameters and their values (Dict)",
12 "github": "The link to the gitHub repository containing the code for this paper, if available (String)"
13}
@@ Appendix A Example Prompts
Table 5: The example JSON file of the table extraction agent with table classification COT.
@@
Appendix B Example LAG-generated Leaderboards
@@
Figure 5 and 6 illustrate two examples generated by LAG. The papers in the first leaderboard are the latest methods of semi-supervised medical image segmentation on the LA dataset from March to December in 2024. The second leaderboard collects the most recent methods for image quality assessment conducted on the LIVE dataset from February 2022 to November 2024. To ensure that the table content is fully displayed, the "model & size" and "hyperparameters selection" within the experimental settings are presented beneath the paper titles.
@@
First and foremost, when viewed holistically, both leaderboards with 20 entries, whether utilizing qwen2.5-14B or GPT-4o as the construction model, exhibit a notably high level of completeness. Upon specific analysis of the missing information, in Leaderboard 1, LAG failed to successfully extract the HD value from "Self-Paced Sample Selection for Barely-Supervised Medical Image Segmentation" (No. 11) because the metric was referred to as 95HD in the original text. Although our design accounts for such situations: our designed agent is required to extract metrics from both text and tables to avoid confusion caused by abbreviations. This design has successfully resolved most of the issues arising from abbreviations, but such errors still occur with a small probability. The absence of metrics in entries No. 13 and No. 20 is acceptable because the original text indeed lacks these metrics. The situation in Leaderboard 2 is similar; the only two missing items (No. 4 and No. 14) are also due to the absence of corresponding results in the original texts.
@@
The higher missing rate in the 5-row leaderboard compared to the 20-row leaderboard for the LIVE dataset can be attributed to the following reasons: When only 5 papers are included, LAG extracts a larger number of metrics, including RMSE, mIoU, and mAcc. The missing values for these metrics are tolerable in a 5-row leaderboard. However, when expanding to a 20-row leaderboard, the excessive number of missing values forces LAG to discard these metrics to ensure that the leaderboard conveys meaningful information.
@@
Secondly, regarding the experimental settings, we observe that in Leaderboard 1, the information on "model & size", "hyperparameters", and "training strategy" is both accurate and comprehensive. Notably, there is a consistent thread throughout the hyperparameters: the portion of labeled data. In contrast, Leaderboard 2 discards the hyperparameter information compared to Leaderboard 3. This is because we require LAG to extract hyperparameter information in a way that not only maintains completeness but also focuses on the intrinsic connections between different items. If the deviation is too large (i.e., if it cannot provide users with a concise and effective summary), the information should be discarded. Therefore, when the number of input papers for LAG increases from 5 to 20, the hyperparameter settings in the topic of image quality assessment do not have a clear and unified theme and thus are ultimately ignored.
Appendix C Cost Analysis

### v2

@@ Appendix B The Use of Large Language Models
Figure 4: Left: Impact of Iteration on League Performance. Right: Pearson Correlation Coefficient values given by four LLMs and human experts. Note that the Pearson Correlation Coefficient is between -1 and 1, the larger value indicates more positive correlations.
Table 5: Results of leaderboard quality generated by leveraging the official APIs provided by OpenReview (openreview-py) and the ACL Anthology (acl-anthology), we systematically harvested camera-ready PDFs from top-tier venues (NeurIPS, ICML, ICLR, ACL, EMNLP, NAACL).
Items Topic-related Quality Model Speed/s Content Quality Recall Precision Coverage Latest Structure Multiaspect 5 $82.36_{\pm 9.42}$ $80.77_{\pm 10.05}$ Qwen2.5-7B $176.41$ $3.96_{\pm 0.42}$ $3.83_{\pm 0.37}$ $3.54_{\pm 0.40}$ $3.78$ Qwen2.5-14B $188.17$ $4.34_{\pm 0.42}$ $4.28_{\pm 0.35}$ $3.95_{\pm 0.31}$ $4.19$ GPT-4o $103.32$ $4.66_{\pm 0.40}$ $83_{\pm 0.34} 4.4$$43_{\pm 0.37}$ $4.64$ O1-preview $117.22$ $4.70_{\pm 0.42}$ $4.85_{\pm 0.65}$ $4.51_{\pm 0.31}$ $4.69$ Human Writing $433$ 4.83 4.93 4.90 4.89 10 $77.25_{\pm 8.17}$ $79.96_{\pm 8.66}$ Qwen2.5-7B $199.76$ $3.63_{\pm 0.47}$ $3.50_{\pm 0.55}$ $4.09_{\pm 0.47}$ $3.74$ Qwen2.5-14B $207.39$ $4.12_{\pm 0.45}$ $3.83_{\pm 0.46}$ $_{\pm 0.32} 3.973.97$ GPT-4o $113.43$ $4.72_{\pm 0.41}$ $4.53_{\pm 0.35}$ $4.72_{\pm 0.35}$ $4.66$ O1-preview $136.57$ $4.61_{\pm 0.32}$ $66_{\pm 0.66} 4.4$$40_{\pm 0.41}$ $4.56$ Human Writing $780$ $4.82$ $4.67$ $4.70$ $4.73$ 15 $70.21_{\pm 7.54}$ $73.90_{\pm 6.83}$ Qwen2.5-7B $243.11$ $3.36_{\pm 0.18}$ $3.29_{\pm 0.22}$ $45_{\pm 0.33} 3.3$73.37 Qwen2.5-14B $255.64$ $66_{\pm 0.25} 3.3$$37_{\pm 0.20}$ $3.55_{\pm 0.27}$ $3.53$ GPT-4o $124.55$ $4.59_{\pm 0.24}$ $4.33_{\pm 0.29}$ $35_{\pm 0.20} 4.4$24.42 O1-preview $137.80$ $4.35_{\pm 0.44}$ $4.16_{\pm 0.23}$ $4.30_{\pm 0.29}$ $4.27$ Human Writing $1176$ $4.70$ $4.53$ $4.51$ $4.58$
@@
Table 6: The prompts of the table extraction (w and w/o the table classification COT procedure) differs in the [EXAMPLE JSON]. It is detailed in the Table 8.
<instruction> You are an expert in summarizing and extracting key content from LaTeX-formatted academic papers on computers and artificial intelligence. Please output your reply in the following JSON format: <format>[EXAMPLE JSON]</format> Key points: In the “selected table’s core results”, other models’ results are of no concern and should be omitted. The table’s header metrics should be the same as the evaluation metrics chosen. The number of items in the “classification of tables” dict should be equal to the “number of tables” int value. These two items help you to identify the main result tables better. All three items about the settings in the JSON output should correspond to the proposed method’s best performance in the selected table. Sometimes in the selected table, the proposed method’s performance may not be unique (e.g., different hyperparameters or training strategies); you need to choose the best one, which usually appears in the last row of the table. If there are multiple tables that meet the requirements (both being the main result table and based on the specified dataset), choose the one with richer information. Workflow example: First, I provide you with an article: <article>[EXAMPLE ARTICLE]</article> Then I specify the dataset as [EXAMPLE DATASET], and you should output: <format>[EXAMPLE RESPONSE]</format> </instruction>
Table 7: Prompt of the leaderboard construction.
<instruction> You are an expert in constructing the Artificial Intelligence leaderboard. Please refer to the content I provide you to answer the user’s questions. The contents I provide you are a number of structured summaries extracted from computer/artificial intelligence papers. You need to build a markdown format leaderboard (showcase the performance of the models on the same dataset, each line representing a specific model) based on the titles, experimental settings, and evaluation metrics of these articles. Please output your reply in the Markdown format. Here, I list a complete example of the question and the answer to help you understand your task. For example, I provide you a list of JSON files containing the extracted content of the articles: [JSON LIST] The expected leaderboard that you generate should be: [EXAMPLE LEADERBOARD]   Pay attention   The leaderboard should be in the Markdown format and reflect all the articles provided! The leaderboard in the dictionary format is forbidden! In the above case, selecting Pre and Rec as the metrics in the final leaderboard is not appropriate because in most articles the corresponding performance values are absent. Here, the target list of extracted content of the articles is as follows: [TARGET JSON LIST] Warning: I need a well-organized markdown-format leaderboard containing all the articles’ information. The leaderboard’s max serial number in the "No." column should equal to the number of articles provided. When selecting metrics, you need to consider their text descriptions. The same metric may have multiple different abbreviations. In the final table, there must not be any duplicate metrics (it is unacceptable to have duplicates where different abbreviations represent the same meaning). Large-scale omissions are not allowed! For each model, only a small portion of the results are missing under the selected metrics. The vast majority of the metrics have corresponding values. The abbreviations for the same metric may be different, but you need to avoid being misled by the abbreviations. Use approximate intersections to select metrics from the given articles, while avoiding a large amount of data waste. Allow some models to have a certain degree of data missing under the selected metrics. The content in the "Experimental Setting" column should be concise and non-descriptive, just a few words. When different articles use different units for the same metric, please note that you need to convert them when integrating them so that the units in the final leaderboard are consistent. For example, 50% is equal to 0.5. "50" and "0.5" should not be presented in the same column of a leaderboard. Check each column corresponding to the selected metrics in the final leaderboard. If more than 60% of the values in that column are missing or represented by placeholders, the metric should be discarded. </instruction>
Table 8: The example JSON file of the table extraction with table classification COT.
1{ 2 "title": "The title of the paper (String)", 3 "number of tables": "The number of tables in the paper, denoted as n (Int)", 4 "classification of tables": { 5 "0": "main-result/comparison", 6 "1": "ablation", 7 "2": "hyper-parameter", 8 "3": "others" 9 }, 10 "selected table’s index": "The index of the main result table focused on the specified dataset [SPECIFIED DATASET], denoted as i (Int)", 11 "metrics": "The evaluation metrics chosen to assess performance of the method proposed in this paper. This information is extracted from the textual portion of the ’Experimental’ related section (String)", 12 "selected table’s metrics": "Metrics used in the selected main result table, it should be almost the same as the metrics extracted from the textual. Remove the latex format syntax (String)", 13 "selected table’s core results": "A dictionary only containing this paper’s model best performance on the selected dataset, with the metrics as keys and the corresponding values (Dict)", 14 "selected table’s settings (model & size)": "In computer vision, the model usually means the backbone architecture of the network, such as ResNet, ViT, and so on. The size can be omitted if not specified. In NLP, the model and size are usually organized as a string, such as ’LLAMA-7B’, ’GPT-3’, and so on (String)", 15 "selected table’s settings (training strategy)": "Training strategy usually refers to the concepts like: fine-tuning, transfer learning, linear-probing, reinforce learning, one-shot, few-shot, prompt-learning, semi/self supervised and so on (String)", 16 "selected table’s settings (hyperparameter selection)": "The hyperparameters used in the model, such as learning rate, batch size, and so on. You should output a dictionary with the hyperparameters and their values (Dict)", 17 "github": "The link to the gitHub repository containing the code for this paper, if available (String)" 18}
Table 9: Leaderboard Quality Criteria.
Criteria Scores Coverage The ratio of the number of papers used for leaderboard generation to the total number of papers searched. $(P_{used}/P_{total})*5$ Latest The ratio of the number of papers published after the certain date to the total number of papers searched. $(P_{new}/P_{total})*5$ Structure Score 1: The structure of the leaderboard lacks logic, making it difficult to understand and navigate. The table header and each row are not clearly organized and connected. Score 2: The structure of the leaderboard have some contents arranged in a disordered or unreasonable manner. However, the overall structure is reasonable and coherent. Score 3: The survey is generally comprehensive in coverage but still misses a few key points that are not fully discussed. Score 4: The structure of the leaderboard is generally reasonably logical, with most header items arranged orderly, though some header items may be repeated or redundant. Score 5: The structure of the leaderboard has good logical consistency, with each line strictly related to the header items and the previous line. But it can be optimized in terms of easy understanding. Multi-Aspect The evaluation metric for multi-leaderboard. Specifically, a research topic $T$ may have $N$ different datasets, and thus we can get $N$ leaderboards, the score of the Multi-Aspect is computed based on the average of all the $N$ scores. $(N_{Coverage}+N_{Latest}+N_{Structure})/(3*N)$.
Appendix C Example Prompts
The prompts of instructing LLMs in different stages of League are illustrated in Box 6, and 7.
@@ Appendix C Example Prompts
Figure 8: The workflow of League from a single table’s perspective.
Table 10: Illustration of the table’s cell entity recognition, where [Uncaptioned image] stands for datasets, [Uncaptioned image] for methods, [Uncaptioned image] for metrics, and [Uncaptioned image] for experimental settings.
[…]
Appendix D Implemenation Details
Table Classification and Table NER
To illustrate the effectiveness of stages 2 and 3. We manually annotated 354 tables from 72 papers, with 197 main results tables, 54 ablation study tables, and 103 others. For Table NER, we follow the annotation criteria of SciIE  (Li et al., 2023) and annotated 336 entities from 28 tables. The entities contain 172 methods, 31 datasets, 49 settings, and 84 metrics. As illustrated in Table 3, we using different prompt strategies: with 0-shot and 1-shot. For prompt with 0-shot, we just provide a brief definition of the sub-categories. For 1-shot, we give an example table for each category besides the definition.
Leaderboard Construction
For proprietary models, we employ official APIs to interact with exclusive LLMs, and the prompts are well-defined. we set temperature = 0.3 and other parameters as default. For open-source models, all experiments are conducted on a single A100 GPU. The input length is set to 128K tokens and the max output tokens is 4096.
Appendix E Example League-generated Leaderboards
@@
Figure 5 and 6 illustrate two examples generated by League. The papers in the first leaderboard are the latest methods of semi-supervised medical image segmentation on the LA dataset from March to December in 2024. The second leaderboard collects the most recent methods for image quality assessment conducted on the LIVE dataset from February 2022 to November 2024. To ensure that the table content is fully displayed, the "model & size" and "hyperparameters selection" within the experimental settings are presented beneath the paper titles.
@@
First and foremost, when viewed holistically, both leaderboards with 20 entries, whether utilizing qwen2.5-14B or GPT-4o as the construction model, exhibit a notably high level of completeness. Upon specific analysis of the missing information, in Leaderboard 1, League failed to successfully extract the HD value from "Self-Paced Sample Selection for Barely-Supervised Medical Image Segmentation" (No. 11) because the metric was referred to as 95HD in the original text. Although our design accounts for such situations, required to extract metrics from both text and tables to avoid confusion caused by abbreviations. This design has successfully resolved most of the issues arising from abbreviations, but such errors still occur with a small probability. The absence of metrics in entries No. 13 and No. 20 is acceptable because the original text indeed lacks these metrics. The situation in Leaderboard 2 is similar; the only two missing items (No. 4 and No. 14) are also due to the absence of corresponding results in the original texts.
@@
The higher missing rate in the 5-row leaderboard compared to the 20-row leaderboard for the LIVE dataset can be attributed to the following reasons: When only 5 papers are included, League extracts a larger number of metrics, including RMSE, mIoU, and mAcc. The missing values for these metrics are tolerable in a 5-row leaderboard. However, when expanding to a 20-row leaderboard, the excessive number of missing values forces League to discard these metrics to ensure that the leaderboard conveys meaningful information.
@@
Secondly, regarding the experimental settings, we observe that in Leaderboard 1, the information on "model & size", "hyperparameters", and "training strategy" is both accurate and comprehensive. Notably, there is a consistent thread throughout the hyperparameters: the portion of labeled data. In contrast, Leaderboard 2 discards the hyperparameter information compared to Leaderboard 3. This is because we require League to extract hyperparameter information in a way that not only maintains completeness but also focuses on the intrinsic connections between different items. If the deviation is too large (i.e., if it cannot provide users with a concise and effective summary), the information should be discarded. Therefore, when the number of input papers for League increases from 5 to 20, the hyperparameter settings in the topic of image quality assessment do not have a clear and unified theme and thus are ultimately ignored.
Appendix F Workflow Illustration
Figure 8 illustrates the transformation of the main experimental table of the LA dataset in the SDCL  (Song & Wang, 2024) through the entire workflow of League. Initially, the table information is presented in the form of visual features within a PDF file. Subsequently, through the processes of crawling and LaTeX integration, the table is extracted and classified into an independent LaTeX format. Further, the table information is structured by the COT table extraction. Finally, after evaluation and iteration, it becomes an entry in the final leaderboard.
Appendix G Cost Analysis